OpenStudy (anonymous):

The radius of an imploding spherical star is decreasing at a rate of 55 miles per hour. When its radius is 100,000 miles in length, at what rate is its surface area changing?

OpenStudy (amistre64):

we will want the derivative of the surface area of a sphere ....

OpenStudy (anonymous):

The volume of a sphere is V=4/3pir^3 so I first take the derivative right?

OpenStudy (amistre64):

we dont need volume since the questions has no information regarding volume

OpenStudy (amistre64):

surface area of a sphere is key

OpenStudy (anonymous):

A=4pir^2

OpenStudy (amistre64):

now rate of change of A is the implicit derivative ....

OpenStudy (anonymous):

da/dt=2pi*r*dr/dt

OpenStudy (amistre64):

forgot the 4 :) but yeah

OpenStudy (amistre64):

they give us r and r' sooo the rest is just multiplicaiton

OpenStudy (anonymous):

wait what four...

OpenStudy (amistre64):

.... the 4 that was a part of the equation for the area to start with of course

OpenStudy (anonymous):

I didnt quite understand that ? Is'nt that the equation for the area of a phere?

OpenStudy (amistre64):

the equation is fine A = 4 pi r^2 take the derivative A = 4 * 2pi r * r'

OpenStudy (amistre64):

we cant just ignore the constant ....

OpenStudy (anonymous):

I searched up the wrong formula lol.

OpenStudy (anonymous):

so the derivative would be da/dt=8pirdr/dt

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

da/dt= -44,000,000 miles/hour